four cells each of emf 2V and internal resistance 1ohm are connected in parallel with an external resistance of 6ohm.The current in the external resistance is?
 

Grade:12

2 Answers

Arun
4 years ago
I) As cells are in parallel, the battery has net emf of 1.5 V and its net internal resistance, r is given by 
1/r = 1 + 1 + 1 + 1= 4 => r = 0.25 ohm. 
(ii) total resistance, R = 0.25 + 6 = 6.25 ohm
(iii) current flowing in the external circuit is
I = 2 / R = 2 / 6.25 = 0.32A
 
Khimraj
4 years ago
 
 battery has net emf of 2 V and its net internal resistance, r is given by 
1/r = 1 + 1 + 1 + 1= 4 => r = 0.25 ohm. 
R = 0.25 + 6 = 6.25 ohm
current 
I = 2 / R = 2 / 6.25 = 0.32A